# TimeMosaic: Temporal Heterogeneity Guided Time Series Forecasting via Adaptive Granularity Patch and Segment-wise Decoding

### v4

@@ Appendix A Acknowledgments
This work was supported by the Innovation Funding of Institute of Computing Technology, Chinese Academy of Sciences under Grant No. E461070.
@@ Appendix H Detailed Results
Models
TimeMosaic SimpleTM TimeFilter xPatch PatchMLP DUET PathFormer iTransformer TimeMxier PatchTST DLinear FreTS
@@
Models
TimeMosaic SimpleTM TimeFilter xPatch PatchMLP DUET LightTS iTransformer TimeMxier PatchTST DLinear FreTS
@@
Models
TimeMosaic SimpleTM TimeFilter DUET iTransformer TimeMixer PatchTST DLinear
@@
$ETTh1$
$\rightarrow$
$ETTh2$
96 0.301 0.341 0.510 0.576 0.335 0.374 0.347 0.400 0.304 0.350
@@
$ETTh1$
$\rightarrow$
$ETTm2$
96 0.212 0.302 0.646 0.563 0.236 0.315 0.255 0.357 0.215 0.304
@@
$ETTh2$
$\rightarrow$
$ETTh1$
96 0.471 0.454 1.130 0.777 0.732 0.577 0.689 0.555 0.485 0.465
@@
$ETTh2$
$\rightarrow$
$ETTm2$
96 0.226 0.308 0.646 0.563 0.253 0.329 0.240 0.336 0.226 0.309
@@
$ETTm1$
$\rightarrow$
$ETTh2$
96 0.331 0.379 0.510 0.576 0.353 0.392 0.365 0.415 0.354 0.385
@@
$ETTm1$
$\rightarrow$
$ETTm2$
96 0.179 0.267 0.646 0.563 0.217 0.294 0.221 0.314 0.195 0.271
@@
Dataset
z-score RevIN

### v5

@@ Appendix A Acknowledgments
This research was partly supported by the National Natural Science Foundation of China under Grant Nos. U24A20232 and 62402475.
@@ Appendix H Detailed Results
Models TimeMosaic SimpleTM TimeFilter xPatch PatchMLP DUET PathFormer iTransformer TimeMxier PatchTST DLinear FreTS
@@
Models TimeMosaic SimpleTM TimeFilter xPatch PatchMLP DUET LightTS iTransformer TimeMxier PatchTST DLinear FreTS
@@
$ETTh1$ $\rightarrow$ $ETTh2$ 96 0.301 0.341 0.510 0.576 0.335 0.374 0.347 0.400 0.304 0.350
@@
$ETTh1$ $\rightarrow$ $ETTm2$ 96 0.212 0.302 0.646 0.563 0.236 0.315 0.255 0.357 0.215 0.304
@@
$ETTh2$ $\rightarrow$ $ETTh1$ 96 0.471 0.454 1.130 0.777 0.732 0.577 0.689 0.555 0.485 0.465
@@
$ETTh2$ $\rightarrow$ $ETTm2$ 96 0.226 0.308 0.646 0.563 0.253 0.329 0.240 0.336 0.226 0.309
@@
$ETTm1$ $\rightarrow$ $ETTh2$ 96 0.331 0.379 0.510 0.576 0.353 0.392 0.365 0.415 0.354 0.385
@@
$ETTm1$ $\rightarrow$ $ETTm2$ 96 0.179 0.267 0.646 0.563 0.217 0.294 0.221 0.314 0.195 0.271
@@
Dataset z-score RevIN